OpenStudy (anonymous):

Find all functions f:ℝ→ℝ that satisfy f(x^3+y^3)=(x^2)f(x)+yf(y^2) for all x,y∈ℝ Sorry, I don't like littering these forums with a question every few minutes, but they're haaaaarrrd.

OpenStudy (jamesj):

Where are these problems coming from? Some competition paper?

OpenStudy (anonymous):

Possibly. These questions were compiled for me by a friend, with solutions attached. :P Math club is a cool place.

OpenStudy (anonymous):

It is fair to say that f(x^3)=xf(x^2)-->f(x^2)=xf(x), right?

OpenStudy (jamesj):

Zero function is a solution, f(t) = 0 Other constant functions aren't a solution, f(t) = c, c not zero And straight lines through the origin, f(t) = mt, are solutions because \[ f(x^3 + y^3) = m(x^3 + y^3) = x^2(mx) + y(my^2) = x^2f(x) + yf(y^2) \] The next step is harder.

OpenStudy (asnaseer):

since \(x^2f(x)\) involves terms just involving the variable \(x\), and \(yf(y^2)\) involves terms just involving the variable \(y\), then is it valid to assume:\[f(x^3+y^3)=f(x^3)+f(y^3)\]and:\[f(x^3)=x^2f(x)\]\[f(y^3)=yf(y^2)\]

OpenStudy (jamesj):

And from this we can deduce that \( f(x^2) = xf(x) \).

OpenStudy (anonymous):

f(y^3)=yf(y^2)-->yf(y^2)=y^2f(y)?

OpenStudy (anonymous):

This is pretty much where I got stuck, when I tried solving it earlier. AND I'M NOT LOOKING THROUGH THE SOLUTIONS.

OpenStudy (asnaseer):

therefore:\[f(x)=mx\]is the only solution - isn't it?

OpenStudy (anonymous):

It is the correct answer, but I... don't see the methodical step between f(x^2)=xf(x) and that conclusion. It seems intuitive, but my intuition is pretty bad.

OpenStudy (anonymous):

I mean, I see it's correct working backwards, but how can I prove that's the only possible scenario?

OpenStudy (asnaseer):

if:\[f(x^3)=x^2f(x)\]then f(x) can only contain terms in x such that multiplying it by \(x^2\) terms them into terms involving \(x^3\) - that was my thought process.

OpenStudy (asnaseer):

*turns them into

OpenStudy (anonymous):

There's no non-simple function out there that could possibly have this be true? I get the feeling that there's some f(x)!=ax for which f(x^n)=nf(x^(n-1)) is true for only a few values of n.

OpenStudy (asnaseer):

I see you believe in conspiracy theories badreferences :)

OpenStudy (anonymous):

Haha.

OpenStudy (jamesj):

First notice that from asnaseer's identities above, if you set x equal to y then \[ x^2f(x) = f(x^3) = xf(x^2) \] Hence for all non-zero \( x \), we have that \[ f(x^2) = xf(x) \] Now, for the moment, I share BR's anxiety there may be other functions out there. Proving their non-existence is not entirely trivial.

OpenStudy (asnaseer):

I just tried wolfram for fun and got this: http://www.wolframalpha.com/input/?i=solve+f%28x%5E2%29%3Dx*f%28x%29 I wonder why it presents the solution as:\[f(x)=\frac{cx}{e^2}\]

OpenStudy (anonymous):

Man, I don't know how to prove nonexistence. >.> I haven't even finished abstract algebra yet, and only just got through introductory proofs.

OpenStudy (jamesj):

Ah ... there other functions. For instance: Define \[ f\colon\mathbb{R}\to\mathbb{R} \] by $$ f(x) \;=\; \begin{cases} 2x & \text{if }x\text{ is algebraic,} \\ 3x & \text{if }x\text{ is transcendental.}\end{cases} $$ Both algebraic and transcendental numbers are closed under the operation of squaring, and therefore \( f(x^2) = x\,f(x) \) for all \( x \). In fact, whenever we can partition the reals into algebraically closed sets, we can define f as different linear functions on each of those sets.

OpenStudy (asnaseer):

Wow James - you are a true genius!

OpenStudy (anonymous):

@JamesJ, that's a lovely fact: "In fact, whenever we can partition the reals into algebraically closed sets, we can define f as different linear functions on each of those sets." I'm keeping this in mind. It explains so many special functions now that I think about it.

OpenStudy (anonymous):

\(x=y=0\) yields \(f(0)=0\) setting \(x=0\) gives \(f(y^3)=yf(y^2)\) similarly \(y=0\) gives \(f(x^3)=x^2f(x)\) so\[f(x^3+y^3)=f(x^3)+f(y^3)\]setting \(u=x^3\) and \(v=y^3\) gives\[f(u+v)=f(u)+f(v)\]which is Cauchy-equation with 2 solutions : \(f(x)=0\) and \(f(x)=cx\).